Given a short run cost function as TC=1/2Q2+40Q, find the minimum value of AVC and MC.
1
Expert's answer
2020-10-12T10:29:50-0400

If TC = 1/2Q2 + 40Q, then:

AVC = VC/Q = 0.5Q + 40,

MC = TC'(Q) = Q + 40.

So, the minimum value of both AVC and MC is 40 at Q = 0.